You took out a loan with an effective annual interest rate of 13 percent. What is the equivalent 18-month interest rate on this loan?

  

Note: I don't want the APR, I want the EPR (the actual interest rate charged over 18 months).
Answer #1

Effective annual interest rate = EAR = 13%

No of months for equivalent interest rate = n = 18

Equivalent 18-month interest rate = EPR for 18 months = (1+EAR)n/12 - 1 = (1+13%)18/12 - 1 = (1.13)3/2 - 1 = 1.201206 - 1 = 0.201206 = 20.1206%

Hence Equivalent 18-month interest rate = 20.1206%